OpenStudy (rea__):

Adam can eat 3 8 of a pizza in 1 16 of a minute. Compute the unit rate by setting up a complex fraction and simplifying. A) 1 pizza per minute B) 3 pizzas per minute C) 6 pizzas per minute D) 9 pizzas per minute

OpenStudy (shelbywyatt):

3/8*16/1=the answer, not so sure

OpenStudy (rea__):

i got 61.18 . am im doing something wrong?

OpenStudy (supersmart1001):

the answer is simple

OpenStudy (supersmart1001):

first u have to make the bottom numbers the same do u know how to do that

OpenStudy (phi):

write the ratio of pizzas divided by minutes

OpenStudy (supersmart1001):

yes thats what I ment

OpenStudy (supersmart1001):

thanks for saying what I said again @phi

OpenStudy (shelbywyatt):

so we go 6/16 divided by 1/16

OpenStudy (phi):

in this case ⅜ pizza divided by 1/16 of a minute \[ \frac{\frac{3}{8} \ pizza}{\frac{1}{16} \ min} \]

OpenStudy (rea__):

3?

OpenStudy (shelbywyatt):

is dividing fractions equal to multiplying its reciprocal

OpenStudy (phi):

of course, you can simplify that fraction by multiplying top and bottom by 16/1 or 16 \[ \frac{\frac{3}{8} \cdot 16\ pizza}{\frac{1}{16}\cdot 16 \ min} \]

OpenStudy (phi):

notice the bottom is 1/16 * 16. what number is that ?

OpenStudy (shelbywyatt):

3/8*16/1

OpenStudy (shelbywyatt):

16

OpenStudy (rea__):

16

OpenStudy (shelbywyatt):

?

OpenStudy (phi):

the bottom is \[ \frac{1}{16} \cdot 16\]

OpenStudy (rea__):

i got 18.56

OpenStudy (phi):

no. can you simplify \[ \frac{\frac{3}{8} \cdot 16\ pizza}{\frac{1}{16}\cdot 16 \ min} \] what does the bottom part simplify to ?

OpenStudy (rea__):

1?

OpenStudy (phi):

yes, so your problem is now \[ \frac{\frac{3}{8} \cdot 16\ pizza}{1 \ min} \] what does the top simplify to ?

OpenStudy (rea__):

1?

OpenStudy (phi):

what is \[ \frac{3}{8}\cdot 16\] notice that is the same as doing \[ 3 \cdot \frac{16}{8} \]

OpenStudy (phi):

I would divide 8 into 16

OpenStudy (rea__):

which is 0.5

OpenStudy (rea__):

im so lost

OpenStudy (rea__):

so the answer is 3?

OpenStudy (rea__):

or 1?

OpenStudy (phi):

you are doing the top part of \[ \frac{\frac{3}{8} \cdot 16\ pizza}{1 \ min} \] 3* 16/8 is 3* 2

OpenStudy (rea__):

6

OpenStudy (phi):

if arithmetic is confusing, you should review how order of operations work. yes you get 6 ⅜ * 16 is 6 so you now have \[ \frac{6 \ pizzas}{1\ min} \]

OpenStudy (phi):

In English, you would say "6 pizzas per minute"

OpenStudy (rea__):

oh so the answer is 6

OpenStudy (phi):

not just 6 6 pizzas per minute

OpenStudy (rea__):

thank you